Prev Next

SW&IT Meeting Minutes 2022-06-13

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20220530
LR20220530 - Bec, Matthieu 72.0
LR20220530 - Mechtley, Brandon 40.0
LR20220530 - Eychaner, Glenn 72.0
LR20220530 - Edmeades, David 0.0
LR20220530 - Grenz, Paul 8.0
LR20220530 - Hooper, Stephen 0.0
LR20220530 - Zhang, Yang 40.0
LR20220530 - Zhang, Xianyu 72.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • sw-ops support cal 2022B
  • disable button patch for AO (XZ/BM)
  • lbti-wfs and INDI (PG)
  • IT backlog
    • jconsole on obs-x (GE)
    • bandwidth checker/exerciser (IT)
    • what is missing on the new-obs that is available on the old-obs (sciops for SH)
    • dms.tucson outside firewall? 150.135.245.20:53, 80, 443 (MB for IT)
  • Career Conversations due June 30!

Team reports

MB

  • Last week
    • OC, management
    • dms refemem recorder @5Hz released
    • dms.tucson satellite issue, fixed itself?
    • ctrl-l rm501, highlight pointer fix
  • This week
    • OC chair, ATT, SOUL-comm RIXes
    • APU AO stats scripts review/improvements
    • shark-vis PAE doc review
    • fedora f35+ preps?

YZ

  • Previous:
    • TCS/AOS patch (worked okay without problems).
    • Several misc. support/MODS/LBC issues.
    • LBC code study.
    • TMS work and software tweak.
  • Next:
    • TMS test on Monday.
    • TSM software and wiki work (and some light refactoring).
    • Some MODS/LBC issues and discussions with Olga.
    • LBC code study and work.

PSG

  • Last week
    • Look into several LBTI observing scripts which were not working, provided fixes.
    • Began looking at and testing LBTI SX/DX WFS machines to determine why some processes are at 100% usage.
  • This week
    • Apply fixes to WFS processes
    • Check in on AZCAM progress.

DE

  • Last week:
    • Mountain backups looking good, schedule in place, minimal impact
    • Obs4 disk swap with LD
      • Good learning experience for him
      • OK is happy that we got her data moved back over
    • JS equipment set up, all looking good
    • JS accounts (IT, Slack)
    • Tucson Synology disk swap/RMA
    • DIMM machine poking after power issues
    • Cindy desktop network(?) issue persists
    • Final PEPSI observer got his credentials for ssh.mountain. Grumble that he didn't bother to test it before he was on his way up to the summit, but such is life.
  • This week:
    • Refine requested data transfer script.
      • I feel that we've been here before
      • Copying data to /dev/null on Tucson Synology
      • Getting 6-10 MB/s on a 1GB file
      • Bandwith utilization graphs don't suggest that the link is a problem
    • Working with JS on more software installs

GE

  • Last week:
    • IT 8355 (Aladin internal deadlocks): Changed threading implementation to avoid deadlocks.
    • IT 5815 (LUCI GUI slowdowns): I found at least one possible cause for this, which has been solved, and am implementing a Java Security Policy to prevent Aladin from making extraneous network requests.
  • This week:
    • I will be taking June 20 and July 21 off, and working half days (mornings) most days between those dates.
    • IT 5815 (LUCI GUI slowdowns)
    • IT 8355 (Aladin internal deadlocks)
    • IT 8571, 8529 (Observer Scheduler GUI).

BM

  • Last week
    • AO SW meeting / StopAo discussion
    • Went over dxadsec testing/deployment with Xianyu.
  • This week:
    • PRs:
      • Upgrade stable with bank-A only and disabled fsm_fault_recovery button.
      • Housekeeper, gain ramp: discuss with Xianyu.
      • Housekeeper: simulate switch BCU timeout
      • CorrectModes error checking
    • Look at Alfio's flao_logs script
    • ITs:
      • OffsetXY problems: (RTDB disconnect in AOG meeting notes + IT 8307)
      • bcu timeouts + ocam not live (8553, 8250/3957)
      • aos rtdb disconnect (8397)

XZ

  • Last week
    • MMC SW improment stopAO freeze modification with Guido
    • Went over dxadsec testing/deployment with Brandon
    • AdSecs power loss recovery with Guido
    • AdSecs shutdown and recovery with Guido
    • disable fault_recovery button on AdSec Control GUI with Guido
    • SOUL comm, AO SW meeting, LBTI DX AO support, etc
  • This week
    • morning AdSec during TMS test with Guido
    • DX ASM flat characterization with Guido
    • Housekeeper, gain ramp: discuss with Brandon
Topic revision: r11 - 13 Jun 2022, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback